# Diffusion approximation of the stationary distribution of a two-level single server queue

### v7

@@ 4.3 Expansion of the asymptotic BAR
$(r\mu^{(r)}b_{1}+o(r))\mathbb{P}[L^{(r)}\leq\ell^{(r)}_{0}]-c^{(r)}_{1}\mu^{(r)}\mathbb{P}[L^{(r)}=0]$
$\quad=\alpha^{(r)}_{e}\Big{(}\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)=\ell^{(r)}_{0}]$
$\hskip 43.05542pt-\mu^{(r)}\mathbb{E}^{(r)}_{e}[\widehat{R}^{(r)}_{s-}1(L^{(r)}(0-)=\ell^{(r)}_{0})]-\lambda^{(r)}_{1}\mathbb{E}_{s}^{(r)}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0}]\Big{)}+o(r),$
because $\mathbb{P}^{(r)}_{s}[L^{(r)}(0)\leq\ell^{(r)}_{0}]-\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)<\ell^{(r)}_{0}]=\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)=\ell^{(r)}_{0}]$ by Lemma 4.1 and $\mathbb{P}(R^{(r)}_{e}\geq 1/r)=o(r^{2})$ and $\mathbb{P}(R^{(r)}_{s}\geq 1/r)=o(r^{2})$ by (4.29). This proves (4.35). Similarly, (4.36) is obtained from (4.5) and (4.5), ∎
We will use Lemma 4.6 to determine the limits of $\mathbb{P}[L^{(r)}\leq\ell^{(r)}_{0}]$ and $\mathbb{P}[L^{(r)}>\ell^{(r)}_{0}]$, equivalently, $\varphi^{(r)}_{1}(0)$ and $\varphi^{(r)}_{2}(0)$, as $r\downarrow 0$ under some conditions.
@@
$e^{r\theta_{i}\ell^{(r)}_{0}}=e^{\theta_{i}\ell_{0}}(1+o(\theta_{i}r)),\quad e^{r\theta_{i}(\ell^{(r)}_{0}+1)}=e^{\theta_{i}\ell_{0}}(1+\theta_{i}r+o(\theta_{i}r)),\quad i=1,2,$ (4.37)
$e^{r\theta_{1}\ell^{(r)}_{0}}-e^{r\theta_{2}\ell^{(r)}_{0}}=e^{\theta_{1}\ell_{0}}-e^{\theta_{2}\ell_{0}}+e^{(\theta_{1}\vee\theta_{2})\ell_{0}}(o(\theta_{1}r)+o(\theta_{2}r)),$ (4.38)
$e^{r\theta_{1}(\ell^{(r)}_{0}+1)}-e^{r\theta_{2}(\ell^{(r)}_{0}+1)}$
$\quad=e^{\theta_{1}\ell_{0}}-e^{\theta_{2}\ell_{0}}+r(e^{\theta_{1}\ell_{0}}\theta_{1}-e^{\theta_{2}\ell_{0}}\theta_{2})+e^{(\theta_{1}\vee\theta_{2})\ell_{0}}(o(\theta_{1}r)+o(\theta_{2}r)).$ (4.39)
@@
These lemmas are proved in Appendix A.4.
We first compute $D^{(r)}(\theta,0)$.
$D^{(r)}(\theta,0)$ $=\alpha^{(r)}_{e}(1-e^{r\theta\ell^{(r)}_{0}})\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)=\ell^{(r)}_{0}]$
$\qquad+\alpha^{(r)}_{e}e^{r\theta\ell^{(r)}_{0}}\mathbb{E}^{(r)}_{e}[(1-e^{-\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s-}})1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
$\qquad+\alpha^{(r)}_{e}(e^{r\theta(\ell^{(r)}_{0}+1)}-1)\mathbb{E}^{(r)}_{s}[e^{-\eta^{(r)}_{1}(r\theta)\widehat{R}^{(r)}_{e}}1(L^{(r)}(0)=\ell^{(r)}_{0})]$
$\qquad+\alpha^{(r)}_{e}\mathbb{E}^{(r)}_{s}[(e^{-\eta^{(r)}_{1}(r\theta)\widehat{R}^{(r)}_{e}}-1)1(L^{(r)}(0)=\ell^{(r)}_{0})]$
$=-\alpha^{(r)}_{e}e^{r\theta\ell^{(r)}_{0}}\mathbb{E}^{(r)}_{e}[e^{-\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s-}}1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
$\qquad+\alpha^{(r)}_{e}e^{r\theta\ell^{(r)}_{0}}e^{r\theta}\mathbb{E}^{(r)}_{s}[e^{-\eta^{(r)}_{1}(r\theta)\widehat{R}^{(r)}_{e}}1(L^{(r)}(0)=\ell^{(r)}_{0})].$ (4.44)
Define ${\mathcal{E}}^{(r)}_{1}$ is defined as
${\mathcal{E}}^{(r)}_{1}$ $=2^{-1}\mathbb{E}^{(r)}_{e}[\{(\mu^{(r)}\sigma^{(r)}_{s})^{2}(\mu^{(r)}\widehat{R}^{(r)}_{s-})-(\mu^{(r)}\widehat{R}^{(r)}_{s-})^{2}\}1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
$\quad-2^{-1}\mathbb{E}^{(r)}_{s}[\{((\lambda^{(r)}_{1}\sigma^{(r)}_{e,1})^{2}+2)(\lambda^{(r)}_{1}\widehat{R}^{(r)}_{e})-(\lambda^{(r)}_{1}\widehat{R}^{(r)}_{e})^{2}\}1(L^{(r)}(0)=\ell^{(r)}_{0})].$ (4.45)
@@
$D^{(r)}(\theta,0)=\alpha^{(r)}_{e}e^{\theta\ell_{0}}(-r\theta\Delta^{(r)}_{1}+(r\theta)^{2}{\mathcal{E}}^{(r)}_{1})+o((r\theta)^{2}),$ (4.46)
because $\mathbb{E}^{(r)}_{s}[S_{3}(\widehat{R}^{(r)}_{e})1(L^{(r)}(0)=\ell^{(r)}_{0})]$ is uniformly bounded for $r\in(0,1]$ by Corollary 4.4.
@@
${\mathcal{E}}^{(r)}_{2}$ $=-2^{-1}\mathbb{E}^{(r)}_{e}[(\mu^{(r)}\sigma^{(r)}_{s})^{2}(\mu^{(r)}\widehat{R}^{(r)}_{s-})-(\mu^{(r)}\widehat{R}^{(r)}_{s-})^{2}]1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
$\quad+2^{-1}\mathbb{E}^{(r)}_{s}[\{(\lambda^{(r)}_{2}(\sigma^{(r)}_{e,2})^{2}+2)(\lambda^{(r)}_{2}\widehat{R}^{(r)}_{e})-(\lambda^{(r)}_{2}\widehat{R}^{(r)}_{e})^{2}\}1(L^{(r)}(0)=\ell^{(r)}_{0})].$ (4.47)
@@
$D^{(r)}(0,\theta)=\alpha^{(r)}_{e}(e^{r\theta\ell^{(r)}_{0}}-1)\mathbb{E}^{(r)}_{e}[e^{-\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s})}1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
$\quad+\alpha^{(r)}_{e}\mathbb{E}^{(r)}_{e}[(e^{-\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s}}-1)1(L^{(r)}(0-)=\ell^{(r)}_{0})]+\alpha^{(r)}_{e}(1-e^{r\theta(\ell^{(r)}_{0}+1)})\mathbb{P}^{(r)}_{s}[L^{(r)}(0)=\ell^{(r)}_{0}]$
$\quad+\alpha^{(r)}_{e}e^{r\theta(\ell^{(r)}_{0}+1)}\mathbb{E}^{(r)}_{s}[(1-e^{-\eta^{(r)}_{2}(r\theta)\widehat{R}^{(r)}_{e}})1(L^{(r)}(0)=\ell^{(r)}_{0})]$
$=\alpha^{(r)}_{e}e^{r\theta\ell^{(r)}_{0}}\big{\{}\mathbb{E}^{(r)}_{e}[e^{-\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s}}1(L^{(r)}(0-)=\ell^{(r)}_{0})]-e^{r\theta}\mathbb{E}^{(r)}_{s}[e^{-\eta^{(r)}_{2}(r\theta)\widehat{R}^{(r)}_{e}}1(L^{(r)}(0)=\ell^{(r)}_{0})]\big{\}}$
$=\alpha^{(r)}_{e}e^{\theta\ell_{0}}\big{\{}\mathbb{E}^{(r)}_{e}[1+r\theta\mu^{(r)}\widehat{R}^{(r)}_{s-}-2^{-1}(r\theta)^{2}[(\mu^{(r)}\sigma^{(r)}_{s})^{2}(\mu^{(r)}\widehat{R}^{(r)}_{s-})-(\mu^{(r)}\widehat{R}^{(r)}_{s-})^{2}]1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
@@
$\qquad-2^{-1}(r\theta)^{2}[(\lambda^{(r)}_{2}\sigma^{(r)}_{e,2})^{2}(\lambda^{(r)}_{2}\widehat{R}^{(r)}_{e})-(\lambda^{(r)}_{2}\widehat{R}^{(r)}_{e})^{2}]1(L^{(r)}(0)=\ell^{(r)}_{0})]\big{\}}+o((r\theta)^{2})$
$=\alpha^{(r)}_{e}e^{\theta\ell_{0}}(r\theta\Delta^{(r)}_{2}+(r\theta)^{2}{\mathcal{E}}^{(r)}_{2})+o((r\theta)^{2}).$ (4.48)
@@
$\frac{1}{2}rc_{1}\mu\sigma_{1}^{2}\big{(}\beta_{1}-\theta_{1}\big{)}\psi^{(r)}_{1}(\theta_{1})=c_{1}\mu\mathbb{P}[L^{(r)}=0]-e^{\theta_{1}\ell_{0}}\alpha^{(r)}_{e}(\Delta^{(r)}_{1}-r\theta_{1}{\mathcal{E}}^{(r)}_{1})+\theta_{1}o(r),$
$\quad=c_{1}\mu(1-e^{\theta_{1}\ell_{0}})\mathbb{P}[L^{(r)}=0]+re^{\theta_{1}\ell_{0}}\big{(}b_{1}\mu\varphi^{(r)}_{1}(0)+\theta_{1}\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}\big{)}+\theta_{1}o(r),$ (4.51)
@@
$\frac{1}{2}rc_{2}\mu\sigma_{2}^{2}\big{(}\beta_{2}-\theta_{2}\big{)}\psi^{(r)}_{2}(\theta_{2})=\alpha^{(r)}_{e}e^{\theta_{2}\ell_{0}}\big{(}\Delta^{(r)}_{2}+r\theta_{2}{\mathcal{E}}^{(r)}_{2}\big{)}+\theta_{2}o(r)$
$\quad=re^{\theta_{2}\ell_{0}}\big{(}\mu b_{2}\varphi^{(r)}_{2}(0)+\theta_{2}\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{2}\big{)}+\theta_{2}o(r).$ (4.52)
@@
$\mathbb{P}[L^{(r)}=0]=\begin{cases}\frac{r}{2c_{1}\mu\ell_{0}}(c_{1}\mu\sigma_{1}^{2}\varphi^{(r)}_{1}(0)+2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1})+o(r),&b=0,\\ \frac{re^{\beta_{1}\ell_{0}}}{c_{1}\mu(e^{\beta_{1}\ell_{0}}-1)}\big{(}b_{1}\mu\varphi^{(r)}_{1}(0)+\beta_{1}\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}\big{)}+o(r),&b\not=0.\end{cases}$ (4.53)
@@
$\mathbb{P}[L^{(r)}=0]$ $=r\frac{c_{1}\mu\sigma_{1}^{2}\big{(}\theta_{1}-\beta_{1}\big{)}\psi^{(r)}_{1}(\theta_{1})+2e^{\theta_{1}\ell_{0}}(b_{1}\mu\varphi^{(r)}_{1}(0)+\theta_{1}\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1})}{2c_{1}\mu(e^{\theta_{1}\ell_{0}}-1)}+o(r).$ (4.54)
@@
$\mathbb{P}[L^{(r)}=0]$ $=r\lim_{\theta_{1}\to 0}\frac{\theta_{1}(c_{1}\mu\sigma_{1}^{2}\psi^{(r)}_{1}(\theta_{1})+2e^{\theta_{1}\ell_{0}}\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1})}{2c_{1}\mu(e^{\theta_{1}\ell_{0}}-1)}+o(r)$
$=\frac{r}{2c_{1}\mu\ell_{0}}(c_{1}\mu\sigma_{1}^{2}\varphi^{(r)}_{1}(0)+2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1})+o(r),$
@@
$\mathbb{P}[L^{(r)}=0]=\frac{re^{\beta_{1}\ell_{0}}}{c_{1}\mu(e^{\beta_{1}\ell_{0}}-1)}\big{(}b_{1}\mu\varphi^{(r)}_{1}(0)+\beta_{1}\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}\big{)}+o(r),$
@@
$\alpha^{(r)}_{e}\Delta^{(r)}_{1}=\begin{cases}\frac{r}{2\ell_{0}}(c_{1}\mu\sigma_{1}^{2}\varphi^{(r)}_{1}(0)+2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1})+o(r),&b_{1}=0,\\ \frac{r}{e^{\beta_{1}\ell_{0}}-1}\Big{(}b_{1}\mu\varphi^{(r)}_{1}(0)+\frac{1}{c_{1}\mu}\beta_{1}e^{\beta_{1}\ell_{0}}\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}\Big{)}+o(r),&b\not=0.\end{cases}$ (4.55)
@@
$\psi^{(r)}_{1}(\theta_{1})$ =$\begin{cases}\frac{e^{\theta_{1}\ell_{0}}-1}{\ell_{0}\theta_{1}}\Big{(}\varphi^{(r)}_{1}(0)+\frac{2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}}{c_{1}\mu\sigma_{1}^{2}}\Big{)}-\frac{2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}}{c_{1}\mu\sigma_{1}^{2}}e^{\theta_{1}\ell_{0}}+o(1),&b_{1}=0,\\ \frac{\beta_{1}}{e^{\beta_{1}\ell_{0}}-1}\frac{e^{\beta_{1}\ell_{0}}-e^{\theta_{1}\ell_{0}}}{\beta_{1}-\theta_{1}}\Big{(}\varphi^{(r)}_{1}(0)+\frac{2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}}{c_{1}\mu\sigma_{1}^{2}}\Big{)}-\frac{2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}}{c_{1}\mu\sigma_{1}^{2}}e^{\theta_{1}\ell_{0}}+o(1),&b_{1}\not=0,\end{cases}$ (4.56)
$\psi^{(r)}_{2}(\theta_{2})$ $=\frac{\beta_{2}e^{\theta_{2}\ell_{0}}}{\beta_{2}-\theta_{2}}\Big{(}\varphi^{(r)}_{2}(0)+\frac{2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{2}}{c_{2}\mu\sigma_{2}^{2}}\Big{)}-\frac{2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{2}}{c_{2}\mu\sigma_{2}^{2}}e^{\theta_{2}\ell_{0}}+o(1).$ (4.57)

### v8

@@ 4.3 Expansion of the asymptotic BAR
$(r\mu^{(r)}b_{1}+o(r))\mathbb{P}[L^{(r)}\leq\ell^{(r)}_{1}]-c^{(r)}_{1}\mu^{(r)}\mathbb{P}[L^{(r)}=0]$
$\quad=\alpha^{(r)}_{e}\Big{(}\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)=\ell^{(r)}_{1}]$
$\hskip 43.05542pt-\mu^{(r)}\mathbb{E}^{(r)}_{e}[\widehat{R}^{(r)}_{s-}1(L^{(r)}(0-)=\ell^{(r)}_{1})]-\lambda^{(r)}_{1}\mathbb{E}_{s}^{(r)}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{1}]\Big{)}+o(r),$
because $\mathbb{P}^{(r)}_{s}[L^{(r)}(0)\leq\ell^{(r)}_{1}]-\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)<\ell^{(r)}_{1}]=\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)=\ell^{(r)}_{1}]$ by Lemma 4.1 and $\mathbb{P}(R^{(r)}_{e}\geq 1/r)=o(r^{2})$ and $\mathbb{P}(R^{(r)}_{s}\geq 1/r)=o(r^{2})$ by (4.30). This proves (4.36). Similarly, (4.37) is obtained from (4.5) and (4.5), ∎
We will use Lemma 4.6 to determine the limits of $\mathbb{P}[L^{(r)}\leq\ell^{(r)}_{1}]$ and $\mathbb{P}[L^{(r)}>\ell^{(r)}_{1}]$, equivalently, $\varphi^{(r)}_{1}(0)$ and $\varphi^{(r)}_{2}(0)$, as $r\downarrow 0$ under some conditions.
@@
$e^{r\theta_{i}\ell^{(r)}_{1}}=e^{\theta_{i}\ell_{1}}(1+o(\theta_{i}r)),\quad e^{r\theta_{i}(\ell^{(r)}_{1}+1)}=e^{\theta_{i}\ell_{1}}(1+\theta_{i}r+o(\theta_{i}r)),\quad i=1,2,$ (4.38)
$e^{r\theta_{1}\ell^{(r)}_{1}}-e^{r\theta_{2}\ell^{(r)}_{1}}=e^{\theta_{1}\ell_{1}}-e^{\theta_{2}\ell_{1}}+e^{(\theta_{1}\vee\theta_{2})\ell_{1}}(o(\theta_{1}r)+o(\theta_{2}r)),$ (4.39)
$e^{r\theta_{1}(\ell^{(r)}_{1}+1)}-e^{r\theta_{2}(\ell^{(r)}_{1}+1)}$
$\quad=e^{\theta_{1}\ell_{1}}-e^{\theta_{2}\ell_{1}}+r(e^{\theta_{1}\ell_{1}}\theta_{1}-e^{\theta_{2}\ell_{1}}\theta_{2})+e^{(\theta_{1}\vee\theta_{2})\ell_{1}}(o(\theta_{1}r)+o(\theta_{2}r)).$ (4.40)
@@
These lemmas are proved in Appendix A.5. We first compute $D^{(r)}(\theta,0)$.
$D^{(r)}(\theta,0)$ $=\alpha^{(r)}_{e}(1-e^{r\theta\ell^{(r)}_{1}})\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)=\ell^{(r)}_{1}]$
$\qquad+\alpha^{(r)}_{e}e^{r\theta\ell^{(r)}_{1}}\mathbb{E}^{(r)}_{e}[(1-e^{-\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s-}})1(L^{(r)}(0-)=\ell^{(r)}_{1})]$
$\qquad+\alpha^{(r)}_{e}(e^{r\theta(\ell^{(r)}_{1}+1)}-1)\mathbb{E}^{(r)}_{s}[e^{-\eta^{(r)}_{1}(r\theta)\widehat{R}^{(r)}_{e}}1(L^{(r)}(0)=\ell^{(r)}_{1})]$
$\qquad+\alpha^{(r)}_{e}\mathbb{E}^{(r)}_{s}[(e^{-\eta^{(r)}_{1}(r\theta)\widehat{R}^{(r)}_{e}}-1)1(L^{(r)}(0)=\ell^{(r)}_{1})]$
$=-\alpha^{(r)}_{e}e^{r\theta\ell^{(r)}_{1}}\mathbb{E}^{(r)}_{e}[e^{-\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s-}}1(L^{(r)}(0-)=\ell^{(r)}_{1})]$
$\qquad+\alpha^{(r)}_{e}e^{r\theta\ell^{(r)}_{1}}e^{r\theta}\mathbb{E}^{(r)}_{s}[e^{-\eta^{(r)}_{1}(r\theta)\widehat{R}^{(r)}_{e}}1(L^{(r)}(0)=\ell^{(r)}_{1})].$ (4.45)
Define ${\mathcal{E}}^{(r)}_{1}$ as
${\mathcal{E}}^{(r)}_{1}$ $=2^{-1}\mathbb{E}^{(r)}_{e}[\{(\mu^{(r)}\sigma^{(r)}_{s})^{2}(\mu^{(r)}\widehat{R}^{(r)}_{s-})-(\mu^{(r)}\widehat{R}^{(r)}_{s-})^{2}\}1(L^{(r)}(0-)=\ell^{(r)}_{1})]$
$\quad-2^{-1}\mathbb{E}^{(r)}_{s}[\{((\lambda^{(r)}_{1}\sigma^{(r)}_{e,1})^{2}+2)(\lambda^{(r)}_{1}\widehat{R}^{(r)}_{e})-(\lambda^{(r)}_{1}\widehat{R}^{(r)}_{e})^{2}\}1(L^{(r)}(0)=\ell^{(r)}_{1})].$ (4.46)
@@
$D^{(r)}(\theta,0)=\alpha^{(r)}_{e}e^{\theta\ell_{1}}(-r\theta\Delta^{(r)}_{1}+(r\theta)^{2}{\mathcal{E}}^{(r)}_{1})+o((r\theta)^{2}),$ (4.47)
because $\mathbb{E}^{(r)}_{s}[S_{3}(\widehat{R}^{(r)}_{e})1(L^{(r)}(0)=\ell^{(r)}_{1})]$ is uniformly bounded for $r\in(0,1]$ by Corollary 4.4.
@@
${\mathcal{E}}^{(r)}_{2}$ $=-2^{-1}\mathbb{E}^{(r)}_{e}[(\mu^{(r)}\sigma^{(r)}_{s})^{2}(\mu^{(r)}\widehat{R}^{(r)}_{s-})-(\mu^{(r)}\widehat{R}^{(r)}_{s-})^{2}]1(L^{(r)}(0-)=\ell^{(r)}_{1})]$
$\quad+2^{-1}\mathbb{E}^{(r)}_{s}[\{(\lambda^{(r)}_{2}(\sigma^{(r)}_{e,2})^{2}+2)(\lambda^{(r)}_{2}\widehat{R}^{(r)}_{e})-(\lambda^{(r)}_{2}\widehat{R}^{(r)}_{e})^{2}\}1(L^{(r)}(0)=\ell^{(r)}_{1})].$ (4.48)
@@
$D^{(r)}(0,\theta)=\alpha^{(r)}_{e}(e^{r\theta\ell^{(r)}_{1}}-1)\mathbb{E}^{(r)}_{e}[e^{-\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s})}1(L^{(r)}(0-)=\ell^{(r)}_{1})]$
$\quad+\alpha^{(r)}_{e}\mathbb{E}^{(r)}_{e}[(e^{-\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s}}-1)1(L^{(r)}(0-)=\ell^{(r)}_{1})]+\alpha^{(r)}_{e}(1-e^{r\theta(\ell^{(r)}_{1}+1)})\mathbb{P}^{(r)}_{s}[L^{(r)}(0)=\ell^{(r)}_{1}]$
$\quad+\alpha^{(r)}_{e}e^{r\theta(\ell^{(r)}_{1}+1)}\mathbb{E}^{(r)}_{s}[(1-e^{-\eta^{(r)}_{2}(r\theta)\widehat{R}^{(r)}_{e}})1(L^{(r)}(0)=\ell^{(r)}_{1})]$
$=\alpha^{(r)}_{e}e^{r\theta\ell^{(r)}_{1}}\big{\{}\mathbb{E}^{(r)}_{e}[e^{-\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s}}1(L^{(r)}(0-)=\ell^{(r)}_{1})]-e^{r\theta}\mathbb{E}^{(r)}_{s}[e^{-\eta^{(r)}_{2}(r\theta)\widehat{R}^{(r)}_{e}}1(L^{(r)}(0)=\ell^{(r)}_{1})]\big{\}}$
$=\alpha^{(r)}_{e}e^{\theta\ell_{1}}\big{\{}\mathbb{E}^{(r)}_{e}[1+r\theta\mu^{(r)}\widehat{R}^{(r)}_{s-}$
$\qquad-2^{-1}(r\theta)^{2}[(\mu^{(r)}\sigma^{(r)}_{s})^{2}(\mu^{(r)}\widehat{R}^{(r)}_{s-})-(\mu^{(r)}\widehat{R}^{(r)}_{s-})^{2}]1(L^{(r)}(0-)=\ell^{(r)}_{1})]$
@@
$\qquad-2^{-1}(r\theta)^{2}[(\lambda^{(r)}_{2}\sigma^{(r)}_{e,2})^{2}(\lambda^{(r)}_{2}\widehat{R}^{(r)}_{e})-(\lambda^{(r)}_{2}\widehat{R}^{(r)}_{e})^{2}]1(L^{(r)}(0)=\ell^{(r)}_{1})]\big{\}}+o((r\theta)^{2})$
$=\alpha^{(r)}_{e}e^{\theta\ell_{1}}(r\theta\Delta^{(r)}_{2}+(r\theta)^{2}{\mathcal{E}}^{(r)}_{2})+o((r\theta)^{2}).$ (4.49)
@@
$\frac{1}{2}rc_{1}\mu\sigma_{1}^{2}\big{(}\beta_{1}-\theta_{1}\big{)}\psi^{(r)}_{1}(\theta_{1})=c_{1}\mu\mathbb{P}[L^{(r)}=0]-e^{\theta_{1}\ell_{1}}\alpha^{(r)}_{e}(\Delta^{(r)}_{1}-r\theta_{1}{\mathcal{E}}^{(r)}_{1})+\theta_{1}o(r),$
$\quad=c_{1}\mu(1-e^{\theta_{1}\ell_{1}})\mathbb{P}[L^{(r)}=0]+re^{\theta_{1}\ell_{1}}\big{(}b_{1}\mu\varphi^{(r)}_{1}(0)+\theta_{1}\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}\big{)}+\theta_{1}o(r),$ (4.52)
@@
$\frac{1}{2}rc_{2}\mu\sigma_{2}^{2}\big{(}\beta_{2}-\theta_{2}\big{)}\psi^{(r)}_{2}(\theta_{2})=\alpha^{(r)}_{e}e^{\theta_{2}\ell_{1}}\big{(}\Delta^{(r)}_{2}+r\theta_{2}{\mathcal{E}}^{(r)}_{2}\big{)}+\theta_{2}o(r)$
$\quad=re^{\theta_{2}\ell_{1}}\big{(}\mu b_{2}\varphi^{(r)}_{2}(0)+\theta_{2}\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{2}\big{)}+\theta_{2}o(r).$ (4.53)
@@
$\mathbb{P}[L^{(r)}=0]=\begin{cases}\frac{r}{2c_{1}\mu\ell_{1}}(c_{1}\mu\sigma_{1}^{2}\varphi^{(r)}_{1}(0)+2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1})+o(r),&b=0,\\ \frac{re^{\beta_{1}\ell_{1}}}{c_{1}\mu(e^{\beta_{1}\ell_{1}}-1)}\big{(}b_{1}\mu\varphi^{(r)}_{1}(0)+\beta_{1}\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}\big{)}+o(r),&b\not=0.\end{cases}$ (4.54)
@@
$\mathbb{P}[L^{(r)}=0]$ $=r\frac{c_{1}\mu\sigma_{1}^{2}\big{(}\theta_{1}-\beta_{1}\big{)}\psi^{(r)}_{1}(\theta_{1})+2e^{\theta_{1}\ell_{1}}(b_{1}\mu\varphi^{(r)}_{1}(0)+\theta_{1}\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1})}{2c_{1}\mu(e^{\theta_{1}\ell_{1}}-1)}+o(r).$ (4.55)
@@
$\mathbb{P}[L^{(r)}=0]$ $=r\lim_{\theta_{1}\to 0}\frac{\theta_{1}(c_{1}\mu\sigma_{1}^{2}\psi^{(r)}_{1}(\theta_{1})+2e^{\theta_{1}\ell_{1}}\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1})}{2c_{1}\mu(e^{\theta_{1}\ell_{1}}-1)}+o(r)$
$=\frac{r}{2c_{1}\mu\ell_{1}}(c_{1}\mu\sigma_{1}^{2}\varphi^{(r)}_{1}(0)+2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1})+o(r),$
@@
$\mathbb{P}[L^{(r)}=0]=\frac{re^{\beta_{1}\ell_{1}}}{c_{1}\mu(e^{\beta_{1}\ell_{1}}-1)}\big{(}b_{1}\mu\varphi^{(r)}_{1}(0)+\beta_{1}\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}\big{)}+o(r),$
@@
$\alpha^{(r)}_{e}\Delta^{(r)}_{1}=\begin{cases}\frac{r}{2\ell_{1}}(c_{1}\mu\sigma_{1}^{2}\varphi^{(r)}_{1}(0)+2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1})+o(r),&b_{1}=0,\\ \frac{r}{e^{\beta_{1}\ell_{1}}-1}\Big{(}b_{1}\mu\varphi^{(r)}_{1}(0)+\frac{1}{c_{1}\mu}\beta_{1}e^{\beta_{1}\ell_{1}}\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}\Big{)}+o(r),&b\not=0.\end{cases}$ (4.56)
@@
$\psi^{(r)}_{1}(\theta_{1})$ =$\begin{cases}\frac{e^{\theta_{1}\ell_{1}}-1}{\ell_{1}\theta_{1}}\Big{(}\varphi^{(r)}_{1}(0)+\frac{2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}}{c_{1}\mu\sigma_{1}^{2}}\Big{)}-\frac{2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}}{c_{1}\mu\sigma_{1}^{2}}e^{\theta_{1}\ell_{1}}+o(1),&b_{1}=0,\\ \frac{\beta_{1}}{e^{\beta_{1}\ell_{1}}-1}\frac{e^{\beta_{1}\ell_{1}}-e^{\theta_{1}\ell_{1}}}{\beta_{1}-\theta_{1}}\Big{(}\varphi^{(r)}_{1}(0)+\frac{2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}}{c_{1}\mu\sigma_{1}^{2}}\Big{)}-\frac{2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}}{c_{1}\mu\sigma_{1}^{2}}e^{\theta_{1}\ell_{1}}+o(1),&b_{1}\not=0,\end{cases}$ (4.57)
$\psi^{(r)}_{2}(\theta_{2})$ $=\frac{\beta_{2}e^{\theta_{2}\ell_{1}}}{\beta_{2}-\theta_{2}}\Big{(}\varphi^{(r)}_{2}(0)+\frac{2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{2}}{c_{2}\mu\sigma_{2}^{2}}\Big{)}-\frac{2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{2}}{c_{2}\mu\sigma_{2}^{2}}e^{\theta_{2}\ell_{1}}+o(1).$ (4.58)